OpenStudy (gudetamas):

Write an equation of the line passing through the point (−10, 3) that is parallel to the line 5x+2y=12.

OpenStudy (anonymous):

Refer to the attachment from Mathematica version 9.

OpenStudy (dayakar):

equation of the line parallel to ax+by+c1=0 is ax+by+ c2=0 (parallel lines varies only constant term means here c1 and k) for finding k we have a point (-10,3)

OpenStudy (dayakar):

equation of the line parallel to 5x+2y=12 is 5x+2y=k--------(1) which passes through (-10,3) substitute this point in (1) 5*(-10) +2*3=k -50+6=k -44=k therefore required equation is 5x+2y=-44